OpenStudy (anonymous):

solve for x, where x is a real number: 12 x^2 = 3x

OpenStudy (anonymous):

12 X^2 - 3X=0 taking 3X common 3X(4X-1)=0 3X=0 & 4X-1=0 X=0 & X=1/4